Yomi:

Find the value of 8w + 6y - 5h, when w = 3, y = 2.5 and h = 3.6 answer choices 21 28 35

supie:

Ok, so first step is to plug in the numbers.

mxddi3:

ok so you have 8w+6y-5h and we are given that w=3, y=2.5 and h=3.6 all we have to do is plug these values in. so 8(3)+6(2.5)-5(3.6)

Yomi:

yea

se07197:

is 21

Yomi:

thx

mxddi3:

do you understand how you got it?

Yomi:

kinda

Yomi:

ill work it out if i have to

mxddi3:

ok. we're here if you need help!

Yomi:

ik

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie Ok, so first step is to plug in the numbers. \(\color{#0cbb34}{\text{End of Quote}}\) So not the equation will change to \[(8)(3)+(6)(2.5)−(5)(3.6)\] Now, \[(8)(3)+(6)(2.5)−(5)(3.6)= blankz?\]

se07197:

cuz u first multiply the numbers and then do adding and the subtract 8(3)=24 6(2.5)=15 5(3.6)=18 24+15=39 39-18= ?